# CheMixHub: Datasets and Benchmarks for Chemical Mixture Property Prediction

### v1

@@ A.1 Dataset licenses
A.2 Details of molecular graph representation
The GNN takes in molecular graphs derived from the SMILES representations of molecules. Each graph, written as $G=(U,V,E)$, consists of a special global vertex $U$ connected to all other vertices $V$, and a set of edges $E$. The global vertex $U$ encodes overall properties of the molecule and is initialized with 200 normalized RDKit descriptors obtained from descriptastorus [29]. The atoms of the molecules are the vertices (nodes), with node vectors $V=\{{v_{i}}\}^{N_{v}}_{i=1}$ for a molecule with $N_{v}$ atoms, where $v_{i}$ are feature vectors encoding atomic properties. Covalent bonds between atoms are represented as edges $E=\{(e_{k},r_{k},s_{k})\}^{N_{e}}_{k=1}$ for a molecule with $N_{e}$ bonds, where $e_{k}$ are feature vectors of edge properties, and $r_{k},s_{k}\in[1,\ldots,N_{v}]$ are indices of the two atoms that the bond joins together. Note $r_{k}\neq s_{k}$, since bonds must be between two different atoms
Table 4: Features for node and edge features of molecular graphs
Node features Categories
Atomic number 1 (hydrogen) to 54 (iodine), UNK
Atom degree 0, 1, 2, 3, 4, 5, UNK
Formal charge -2, -1, 0, 1, 2, UNK
Chirality unspecified, CW, CCW, other, UNK
Number of hydrogens 0, 1, 2, 3, 4, 5, 6, 7, 8, UNK
Hybridization sp, sp2, sp3, sp3d, sp3d2, UNK
Aromatic True/False
Edge features Categories
Bond type single, double, triple, aromatic, UNK
Is conjugated True/False
In ring True/False
Stereo-configuration none, $Z$, $E$, cis, trans, any, UNK
A.3 Compute resources details
A.4 Training details
Each run was performed for 500 epochs using the Adam optimizer [30], with a batch size set to 1024. Early stopping was implemented with patience set to 100. Two different learning rates were used to train the models end-to-end, one for the molecular-level model and one for the rest of the model. The splits used are specified in Section 3.4, further details on hyperparameter tuning can be found in Section A.5.
A.5 Hyperparameter search
For each task, the search was performed using Weights & Biases [5] with the BOHB algorithm [20] and a budget of 160 runs. 80 runs were allocated to the GNN-based molecular representations and 80 to CLMs and descriptors runs. Each run was performed for 500 epochs with early stopping patience set to 100. The search was conducted using the first split of the 5-fold random CV splits (70/10/20 training/validation/test split). The search space is defined as follows
  • General hyper-parameters:
    • Loss type: ["mae", "mse"]
  • Molecular-level hyper-parameters:
    • FiLM layer activation function: ["sigmoid", "relu"]
  • Mixture-level hyper-parameters:
    • Mixture interaction module: ["self attention", "deepset"]
    • MLP head in self-attention: ["True", "False"]
    • Embedding dimension: [32, 64, 96, 128]
    • Number of layers: [0, 1, 2, 3]
    • Number of attention heads: [1, 4, 8, 16]
    • Output dimension: [96, 128, 256]
    • Mixture context aggregation type: ["concatenate", "film"]
  • Predictive head hyper-parameters:
    • Embedding dimension: [64, 128, 192, 256, 320]
    • Number of layers: [1, 2, 3]
  • GNN hyper-parameters:
    • Embedding dimension: [64, 128, 192, 256, 320]
    • Number of layers: [2, 3, 4]
A.6 XGBoost modeling
The XGBoost model was given a maximum of 1,000 estimators and tree depth of 1,000 except for the full NIST task, where a maximum of 250 estimators and a tree depth of 250 was used. To ensure the model does not overfit, we use the validation set for early stopping, with a patience of 25 epochs. The model is trained with mean squared error, with a learning rate of 0.01.
A.7 IlThermo Dataset curation details
@@
We use the IlThermoPy package to retrieve IlThermo entries, selecting entries that are either binary or ternary mixtures and corresponding to our property of choice (for the scope of this paper, we limit ourselves to viscosity and ionic conductivity properties) [26]. We remove mixture that exhibits multiple phases behavior and are not liquid at the indicated temperature. We apply a natural logarithm transformation to the viscosity and ionic conductivity values present in IlThermo to make the range of values easier to learn. We also constrain the pressure range to be near the standard value of 1 atm or 101.325 kPa by applying a ±$\pm 2$kPa threshold on pressure values.
@@
A.8 Additional metrics for performances across tasks
In addition to the MAE results reported in Section 4.2, we report the results compiled from the CV splits for all models evaluated in terms of Pearson correlation coefficient $\rho$ and Kendall ranking coefficient $\tau$ in Table 5 and 6, respectively.
@@ A.8 Additional metrics for performances across tasks
Molecular rep. Mixture rep. Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{mix}$ $\Delta H_{vap}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular rep. Mixture rep. IlThermo MON NIST Olfaction
@@
Molecular rep. Mixture rep. Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{mix}$ $\Delta H_{vap}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular rep. Mixture rep. IlThermo MON NIST Olfaction
@@
A.9 Additional benchmark
Table 7: DiffMix tasks summary. T indicates Temperature dependency. Mole Fractions indicates if mole fractions of the mixture components are available. Arrhenius relationship indicates if the task can be modeled using the Arrhenius equation. Exp. indicates if the data was obtained from wet-lab experiments or simulations.
Tasks Units Datapoints Max # Components # Unique Mixtures # Unique Molecules Mixture Context Mole Fractions Arrhenius Relationship Exp.
@@ A.9 Additional benchmark
Battery electrolytes—mixtures of salts and solvents—have been optimized to facilitate ion transport, prevent electron transfer, and stabilize electrode-electrolyte interfaces to produce energy-dense and durable battery systems [71, 21]. The DiffMix dataset is a collection of three tasks centered around thermodynamic and transport properties predictions of electrolytes originally gathered by Zhu et al. [80]. This data is under the CC BY-NC-ND 4.0 license, and we therefore cannot include it as part of our dataset.
  • Excess molar enthalpy $H^{E}_{m}$: The excess molar enthalpy reflects changes in intermolecular interactions that occur during the mixing of different components [76]. It shows the non-ideality of the final solution and gives an explanation about enthalpic effects [48]. In particular, differences in molecular shape, size, and interaction types between components—along with variations in temperature and pressure—can lead to either an increase or a decrease in excess molar enthalpy [35, 62]. DiffMix dataset includes 631 $H^{E}_{m}$ data points curated from literature, covering 34 unique mixtures composed of 35 organic compounds across varying compositions. We rescaled the original range of the DiffMix excess molar enthalpy task from J/mol to kJ/mol to avoid passing big values to the neural networks.
  • Excess molar volume $V^{E}_{m}$: The excess molar volume represents the deviation from ideal mixing volume. It exhibits a non-linear dependence on mole fraction [10] and temperature [65]—often showing a U-shaped trend with concentration and a decrease in absolute values as temperature increases. At higher temperatures, the dependence may shift to an S-shaped profile, making accurate prediction particularly challenging [70]. DiffMix dataset includes 1069 $V^{E}_{m}$ data points curated from literature, covering 28 unique mixtures composed of 25 organic compounds.
  • Ionic conductivity $\kappa$: The ionic conductivity of the electrolyte is known as a key parameter to evaluate the performance of the solution in practical engineering applications. In the context of batteries, $\kappa$ changes considerably with the change of the electrolyte concentration [77]. DiffMix dataset includes 24,822 mixtures of single-salt-ternary-solvent electrolyte solutions generated using Advanced Electrolyte Model [22], and covering arbitrary combinations of two unique salts and six organic carbonate solvents at different concentration.
@@
Molecular rep. Mixture rep. DiffMix
@@
Molecular rep. Mixture rep. DiffMix
@@
Molecular rep. Mixture rep. DiffMix
@@ A.10 Modeling salts
We constructed a 200-dimensional molecular embedding space using RDKit 2D descriptors obtained from descriptasorus [29], incorporating both salts and fragments for all the tasks in CheMixHub. The number of unique salts is 824, and the number of fragments is 476. This space was projected into two dimensions using UMAP to visualize structural relationships, Figure 5. As shown in the UMAP plot, The resulting plot shows that salts (blue triangles) and fragments (orange circles) broadly co-localize, with many salts embedded near fragment clusters.
To quantify these observations, we computed cosine distances between each salt and the fragment-only descriptor space. The resulting distribution confirms that the vast majority of salts lie within a narrow cosine distance range centered around 0.04–0.05, with very few exceeding 0.1, Figure 6. In RDKit descriptor space, such low distances imply near-identity in structural features.
From these, we can observe that most salts appear to retain descriptor-level similarity to their constituent fragments. However, there is a subset which introduces structural changes significant enough to shift them away from the fragment space.

### v2

@@ A.1 Dataset licenses
A.2 Additional statistics on molecules for each of the 11 tasks in CheMixHub
Table 6: Additional statistics on molecules for each of the 11 tasks in CheMixHub.
Dataset Tasks
Avg. #
Atoms/Mol
Max #
Atoms/Mol
Min #
Atoms/Mol
Avg #
Fragments
Max #
Fragments
Avg
Molecular
Weight
Avg #
Rotatable
Bonds
Avg
Components
Mixture
Miscible solvents $\rho$ 8.28±3.17 18 3 1.0±0.0 1 123.73±43.96 3.40±3.17 3.72±1.08
$\Delta H_{\mathrm{mix}}$
$\Delta H_{\mathrm{vap}}$
IlThermo $\ln(\kappa)$ 15.80±9.28 77 1 1.76±0.54 4 250.91±145.56 5.12±6.13 2.21±0.41
$\ln(\eta)$ 17.33±10.73 62 1 1.85±0.59 4 280.30±174.57 5.76±6.51 2.40±0.49
NIST Viscosity $\ln(\eta_{\mathrm{NIST-full}})$ 12.90±8.98 95 1 1.50±0.70 8 203.98±135.28 4.00±5.60 1.88±0.33
$\ln(\eta_{\mathrm{NIST}})$ 9.12±4.71 63 1 1.0±0.0 1 140.52±73.17 3.14±3.79 1.92±0.28
Drug solubility $\ln(S)$ 14.48±9.16 51 1 1.11±0.33 3 212.40±128.17 2.37±2.45 1.91±0.29
Solid Polymer Electrolyte $\ln(\kappa)$ 30.86±47.75 676 2 1.24±0.44 3 473.36±738.30 18.11±33.19 2.24±0.67
Olfactory mixtures Perceptual similarity 9.53±3.43 21 3 1.0±0.0 1 135.67±45.03 2.72±2.29 13.30±10.51
Fuel mixtures MON 7.93±1.94 12 2 1.0±0.0 1 110.66±26.19 1.71±1.69 5.69±14.24
A.3 IlThermo Dataset curation details
@@
We use the IlThermoPy package to retrieve IlThermo entries, selecting entries that are either binary or ternary mixtures and corresponding to our property of choice (for the scope of this paper, we limit ourselves to viscosity and ionic conductivity properties) [ilthermopy]. We remove mixture that exhibits multiple phases behavior and are not liquid at the indicated temperature. We apply a natural logarithm transformation to the viscosity and ionic conductivity values present in IlThermo to make the range of values easier to learn. We also constrain the pressure range to be near the standard value of 1 atm or 101.325 kPa by applying a $\pm 2$kPa threshold on pressure values.
@@
A.4 Details of molecular graph representation
The GNN takes in molecular graphs derived from the SMILES representations of molecules. Each graph, written as $G=(U,V,E)$, consists of a special global vertex $U$ connected to all other vertices $V$, and a set of edges $E$. The global vertex $U$ encodes overall properties of the molecule and is initialized with 200 normalized RDKit descriptors obtained from descriptastorus [Kelley2024-kd]. The atoms of the molecules are the vertices (nodes), with node vectors $V=\{{v_{i}}\}^{N_{v}}_{i=1}$ for a molecule with $N_{v}$ atoms, where $v_{i}$ are feature vectors encoding atomic properties. Covalent bonds between atoms are represented as edges $E=\{(e_{k},r_{k},s_{k})\}^{N_{e}}_{k=1}$ for a molecule with $N_{e}$ bonds, where $e_{k}$ are feature vectors of edge properties, and $r_{k},s_{k}\in[1,\ldots,N_{v}]$ are indices of the two atoms that the bond joins together. Note $r_{k}\neq s_{k}$, since bonds must be between two different atoms
Table 7: Features for node and edge features of molecular graphs
Node features Categories
Atomic number 1 (hydrogen) to 54 (iodine), UNK
Atom degree 0, 1, 2, 3, 4, 5, UNK
Formal charge -2, -1, 0, 1, 2, UNK
Chirality unspecified, CW, CCW, other, UNK
Number of hydrogens 0, 1, 2, 3, 4, 5, 6, 7, 8, UNK
Hybridization sp, sp2, sp3, sp3d, sp3d2, UNK
Aromatic True/False
Edge features Categories
Bond type single, double, triple, aromatic, UNK
Is conjugated True/False
In ring True/False
Stereo-configuration none, $Z$, $E$, cis, trans, any, UNK
A.5 Compute resources details
A.6 Training details
Each run was performed for 500 epochs using the Adam optimizer [kingma2014adam], with a batch size set to 1024. Early stopping was implemented with patience set to 100. Two different learning rates were used to train the models end-to-end, one for the molecular-level model and one for the rest of the model. The splits used are specified in Section 3.4, further details on hyperparameter tuning can be found in Section A.7.
A.7 Hyperparameter search
For each task, the search was performed using Weights & Biases [wandb] with the BOHB algorithm [falkner2018bohb] and a budget of 160 runs. 80 runs were allocated to the GNN-based molecular representations and 80 to CLMs and descriptors runs. Each run was performed for 500 epochs with early stopping patience set to 100. The search was conducted using the first split of the 5-fold random CV splits (70/10/20 training/validation/test split). The search space is defined as follows
  • General hyper-parameters:
    • Loss type: ["mae", "mse"]
  • Molecular-level hyper-parameters:
    • FiLM layer activation function: ["sigmoid", "relu"]
  • Mixture-level hyper-parameters:
    • Mixture interaction module: ["self attention", "deepset"]
    • MLP head in self-attention: ["True", "False"]
    • Embedding dimension: [32, 64, 96, 128]
    • Number of layers: [0, 1, 2, 3]
    • Number of attention heads: [1, 4, 8, 16]
    • Output dimension: [96, 128, 256]
    • Mixture context aggregation type: ["concatenate", "film"]
  • Predictive head hyper-parameters:
    • Embedding dimension: [64, 128, 192, 256, 320]
    • Number of layers: [1, 2, 3]
  • GNN hyper-parameters:
    • Embedding dimension: [64, 128, 192, 256, 320]
    • Number of layers: [2, 3, 4]
A.8 XGBoost modeling
The XGBoost model was given a maximum of 1,000 estimators and tree depth of 1,000 except for the NIST-full task, where a maximum of 250 estimators and a tree depth of 250 was used. To ensure the model does not overfit, we use the validation set for early stopping, with a patience of 25 epochs. The model is trained with mean squared error, with a learning rate of 0.01.
A.9 Additional metrics for performances across tasks
In addition to the MAE results reported in Section 4.2, we report the results compiled from the CV splits for all models evaluated in terms of Pearson correlation coefficient $\rho$ and Kendall ranking coefficient $\tau$ in Table 8 and 9, respectively.
@@ A.9 Additional metrics for performances across tasks
Molecular
rep.
Mixture
rep.
Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{\mathrm{mix}}$ $\Delta H_{\mathrm{vap}}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular
rep.
Mixture
rep.
IlThermo MON NIST Olfaction
@@
Molecular
rep.
Mixture
rep.
Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{\mathrm{mix}}$ $\Delta H_{\mathrm{vap}}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular
rep.
Mixture
rep.
IlThermo MON NIST Olfaction
@@
A.10 Additional transfer-learning benchmark
We evaluated transfer learning capabilities of two models trained on different datasets and tasks: the best deep learning model trained on the Miscible Solvent $\Delta H_{\mathrm{vap}}$ task and the other one trained on the Motor Octane Number (MON) task (according to Section 4.2). We compare these fine-tuned models to the best performing models for these tasks found in Section 4.2 (see Table 2).
We observe a simple fine-tuning approach of the best Deep Learning models for each task on another task from a different dataset does not yield good performance, especially compared to "in-dataset" finetuning results above, which could suggest the models are overfitting to their respective tasks. An interesting experimental set up to further answer this questions would be to evaluate multi-task learning capabilities of these models across datasets, which should be easily implementable thanks to our unified framework.
Table 10: Transfer learning capabilities of models across the Miscible Solvent (MS) $\Delta H_{\mathrm{vap}}$ task and the MON task. Metrics are reported on 5-fold random CV splits. The mean and standard deviation are reported. The best model statistics are taken from Section 4.2 and Appendix A.9.
Fine-tuning Dataset
Best model
Original Dataset
Pearson $\rho$ ($\uparrow$) MAE ($\downarrow$) Kendall $\tau$ ($\uparrow$)
MON MON 0.913 ± 0.019 4.570 ± 0.348 0.781 ± 0.029
MS-$\Delta H_{\mathrm{vap}}$ 0.160 ± 0.108 33.199 ± 1.606 0.144 ± 0.056
Miscible Solvent $\Delta H_{\mathrm{vap}}$ MS-$\Delta H_{\mathrm{vap}}$ 0.999 ± 0.000 0.071 ± 0.002 0.976 ± 0.001
MON 0.501 ± 0.095 1.582 ± 0.095 0.296 ± 0.067
A.11 Additional benchmark
Table 11: DiffMix tasks summary. T indicates temperature dependency. Mole Fractions indicates mole fractions availability. Arrhenius relationship indicates if the task can be modeled using the Arrhenius equation. Exp. indicates if the data was obtained from wet-lab experiments or simulations.
Tasks
Units
Datapoints
Max #
Components
# Unique
Mixtures
# Unique
Molecules
Mixture
Context
Mole
Fractions
Arrhenius
Relationship
Exp.
@@ A.11 Additional benchmark
Battery electrolytes—mixtures of salts and solvents—have been optimized to facilitate ion transport, prevent electron transfer, and stabilize electrode-electrolyte interfaces to produce energy-dense and durable battery systems [yu2020molecular, fan2019all]. The DiffMix dataset is a collection of three tasks centered around thermodynamic and transport properties predictions of electrolytes originally gathered by Zhu et al. [zhu2024differentiable]. This data is under the CC BY-NC-ND 4.0 license, and we therefore cannot include it as part of our dataset.
  • Excess molar enthalpy $H^{E}_{m}$: The excess molar enthalpy reflects changes in intermolecular interactions that occur during the mixing of different components [zhang2015excess]. It shows the non-ideality of the final solution and gives an explanation about enthalpic effects [qian2013enthalpy]. In particular, differences in molecular shape, size, and interaction types between components—along with variations in temperature and pressure—can lead to either an increase or a decrease in excess molar enthalpy [lifi2020measurement, verdes2024measurement]. DiffMix dataset includes 631 $H^{E}_{m}$ data points curated from literature, covering 34 unique mixtures composed of 35 organic compounds across varying compositions. We rescaled the original range of the DiffMix excess molar enthalpy task from J/mol to kJ/mol to avoid passing big values to the neural networks.
  • Excess molar volume $V^{E}_{m}$: The excess molar volume represents the deviation from ideal mixing volume. It exhibits a non-linear dependence on mole fraction [cade2014comparing] and temperature [wei2020effect]—often showing a U-shaped trend with concentration and a decrease in absolute values as temperature increases. At higher temperatures, the dependence may shift to an S-shaped profile, making accurate prediction particularly challenging [yang2003excess]. DiffMix dataset includes 1069 $V^{E}_{m}$ data points curated from literature, covering 28 unique mixtures composed of 25 organic compounds.
  • Ionic conductivity $\kappa$: The ionic conductivity of the electrolyte is known as a key parameter to evaluate the performance of the solution in practical engineering applications. In the context of batteries, $\kappa$ changes considerably with the change of the electrolyte concentration [zhang2020experimental]. DiffMix dataset includes 24,822 mixtures of single-salt-ternary-solvent electrolyte solutions generated using Advanced Electrolyte Model [gering2017prediction], and covering arbitrary combinations of two unique salts and six organic carbonate solvents at different concentration.
@@
Molecular
rep.
Mixture
rep.
DiffMix
@@
Molecular
rep.
Mixture
rep.
DiffMix
@@
Molecular
rep.
Mixture
rep.
DiffMix
@@ A.12 Modeling salts
We constructed a 200-dimensional molecular embedding space using RDKit 2D descriptors obtained from descriptasorus [Kelley2024-kd], incorporating both salts and fragments for all the tasks in CheMixHub. The number of unique salts is 824, and the number of fragments is 476. This space was projected into two dimensions using UMAP to visualize structural relationships, Figure 5. As shown in the UMAP plot, The resulting plot shows that salts (blue triangles) and fragments (orange circles) broadly co-localize, with many salts embedded near fragment clusters. To quantify these observations, we computed cosine distances between each salt and the fragment-only descriptor space. The resulting distribution confirms that the vast majority of salts lie within a narrow cosine distance range centered around 0.04–0.05, with very few exceeding 0.1, Figure 6. In RDKit descriptor space, such low distances imply near-identity in structural features. From these, we can observe that most salts appear to retain descriptor-level similarity to their constituent fragments. However, there is a subset which introduces structural changes significant enough to shift them away from the fragment space.